# Semantic Regexes: Auto-Interpreting LLM Features with a Structured Language

### v1

@@ 5.3 Semantic Regexes Reflect Feature Complexity
Beyond matching the accuracy of natural language descriptions and offering benefits like conciseness and consistency, the structured format of semantic regexes offers additional affordances for model interpretability. Each semantic regex is built from primitives that span increasing levels of abstraction, where symbols match specific characters, lexemes extend to syntactic variants, and fields encode semantic relationships. Moreover, the number of components in a semantic regex is also a proxy for feature complexity, where features described using a single primitive (e.g., [:symbol left:] ) are typically conceptually simpler than features that require multiple compositions of primitives and modifiers (e.g., @{:context political affiliations:}([:symbol left:]||[:symbol right:]) ).
We use the level of abstraction encoded in semantic regexes to measure feature complexity, finding that features become more complex deeper in the model (Figure 4). We generate semantic regexes for 1,000 features per layer in GPT-2-RES-25k, Gemma-2-2B-RES-16k, and Gemma-2-2B-RES-65k. While early-layer features are described by smaller and simpler semantic regexes, longer and more abstract semantic regexes are needed to describe later-layer features. In particular, we find that the average number of components per semantic regex (i.e., symbols, lexemes, fields, and contexts) steadily increases across layers. This shift is mirrored in the composition of a semantic regex, where early layers have a greater proportion of single-primitive descriptions which decreases in favor or combinations of primitives and modifiers, particularly sequence and alternation compositions. Similarly, we observe that the types of primitives reflect increasing feature complexity. While all semantic regexes contain primitives, we see a decrease in low-level primitives, like lexemes, and an increase in fields (the most abstract primitive). We show an example of increasing feature complexity from early to later layers in Figure 3 bottom.
Together, these trends indicate that later layers require longer and more abstract semantic regexes to capture their complex feature behaviors. This aligns with prior research demonstrating that later layers encode increasingly complex representations (Jin et al., 2025; Tenney et al., 2019; Sun et al., 2025). However, unlike prior methods that rely on model probes or feature testing, we are able to read this complexity directly from the semantic regex feature description. As a result, while like natural language descriptions, semantic regexes allow us understand individual features, their structure also allows us to interpret entire model attributes.
@@ 5.4 Semantic Regexes Help People Build Mental Models of LLM Features
A common role of feature descriptions is to convey the feature’s behavior to a human interpreter (Ameisen et al., 2025; Lin, 2023). Thus, we investigate whether semantic regexes help people form accurate mental models of LLM features. To do so, we conducted a 24-person user study, testing whether AI experts could generate data that did and did not activate a feature given only its description (full details in Appendix B). We selected 12 features from GPT-2-RES-25k that represented diverse activation patterns, spanned model layers, and had accurate feature descriptions. Given a feature description, participants were asked to generate three activating phrases and one near-miss counterfactual. Each participant generated phrases for three natural language and three semantic regex feature descriptions, resulting in 425 positive and 143 counterfactual phrases.
Forming an accurate mental model of an LLM feature means being able to express the feature’s decision boundary, i.e., what does and does not activate the feature. We quantify this by measuring the difference between the feature’s maximum activation on each participant’s positive phrases and their counterfactual phrase. Small or negative values indicate that the participant had difficulty distinguishing activating and non-activating phrases, while large positive values reflect an understanding of the feature’s decision boundary. We compare the mean differences of max-acts and semantic-regex descriptions of the same feature, finding that participants scored higher using semantic regex descriptions on 9 of 12 features (Figure 5).
@@
Although both description types were accurate, natural language often introduced extraneous details that misled participants. For instance, one feature activated on variants of the phrase “expected to”, but its natural language description included “‘expected to’ is frequently used to indicate anticipation or prediction regarding future events or outcomes”. A participant over-indexed on the additional detail, expecting the highly activating phrase “He does not know meaning of the phrase expected to” to be a counterexample because it did not indicate anticipation. In contrast, [:lexeme expect:][:symbol to:] expresses only the activation pattern and enabled all participants to generate strongly activating positives and non-activating counterfactuals.
We also find that semantic regexes reduce ambiguity by conveying activation patterns via example. For a feature that activates on times followed by “p.m.”, the semantic regex specifies this pattern directly as [:field time:][:symbol p.m.:] , and as a result, every participant included “p.m.” in their positive phrases. In contrast, natural language descriptions can leave room for interpretation, where one participant misinterpreted “reference to time, specifically indicating hours in the ‘p.m.’ format” as any time in the evening and left off the critical “p.m.” A similar pattern occurred in a feature for political conjunctions, where the semantic regex directly included [:symbol and:]|[:symbol or:] , making the intended conjunction pattern unambiguous.
While semantic regexes’ concise examples generally benefited participants, there were cases where the verbosity of natural language was more effective. In the two features where natural language substantially outperformed semantic regexes, the semantic regexes were accurate but too minimal to elicit strongly activating phrases from participants. For example, one feature activates on days of the week used within full sentences. The conciseness of the semantic regex [:field days of the week:] led many participants to produce single-word outputs, like “Monday”, which activated slightly. However, the natural language description “days of the week indicating specific events” prompted longer phrases with larger activations.
@@ 6 Conclusion
We introduce semantic regexes, a structured language for automatically describing LLM features. The semantic regex language is grounded in our current understanding of how LLM features operate. Each primitive and modifier is designed to reflect patterns observed in interpretability research — i.e., that features often respond to exact tokens ( symbol ), syntactically related word forms ( lexeme ), and semantically related categories ( field ), and their activations are often shaped by domain constraints ( context ) or co-occurrence with other patterns (Lin, 2023; Templeton et al., 2024). As our understanding of LLM representations grows, we expect the semantic regex language will evolve, with new components added or altered to capture novel feature behaviors. Moreover, just as there are many languages for programming, alternative structured languages for interpretability may be developed with affordances that match particular interpretability goals, such as highlighting input vs. output features, describing particular model components (e.g., attention heads), or exposing safety-relevant aspects of features. The affordances of consistency, conciseness, and abstraction that we build into semantic regexes expose a broader design space of structured languages for interpretability that we hope will improve our understanding and control of LLMs.
@@ Acknowledgments
We thank our colleagues at Apple and our user study participants for their contributions to this research, with a special mention to Hadas Kotek, Lindia Tjuatja, and Mary Beth Kery for giving us feedback on early versions of this work.
Appendix A Additional Results
@@
Appendix B User Study Protocol
Here we provide the protocol for our user study in Section 5.4.
@@ Appendix B User Study Protocol
We recruited 24 participants from within our organization. We selected participants who work in AI, representing practitioners we expect to interpret LLM features without overly constraining to interpretability experts.
@@
We selected 12 features from GPT-2-RES-25k that had both accurate semantic regex and natural language feature descriptions, came from different layers of the model, and represented a diverse set of activation patterns. We show all 12 features and thier descriptions in Figure A2.
@@
Participants were tasked with writing phrases that matched or were a close counterfactual to a given feature description. Each participant was shown three natural language descriptions and three semantic regex descriptions. To avoid carryover effects, participants saw natural language and semantic regex descriptions from different features. To avoid learning effects, we randomized the order, with half the participants seeing natural language descriptions first and the other half seeing semantic regexes first.
Each participant completed the task for three natural language feature descriptions (max-acts) and three semantic-regex feature descriptions. We showed each participant features of varying complexity, with half the participants seeing the feature descriptions in the left column of Figure A2 and the other half seeing feature descriptions from the right column of Figure A2.
The task was completed asynchronously and distributed via a survey link. The survey began with an introduction to the task, the semantic regex language, and an example of matching and counterfactual phrases (Figure A3). Then, participants were shown an explanation of the feature description type they would see first, followed by three feature descriptions they were asked to generate phrases for. This was repeated with the alternate feature description type and the final three explanations (Figures A4 and A5). For each feature description, they were asked to “Write 3 phrases/sentences that match the feature description” and “Write 1 phrase/sentence that is a counterfactual for the semantic regex feature description” (Figures A6 and A7).
@@
Appendix C Implementation Details
C.1 Feature Description Methods

### v2

@@ 5.3 Semantic Regexes Reflect Feature Complexity
Beyond matching the accuracy of natural language descriptions and offering benefits like conciseness and consistency, the structured format of semantic regexes offers additional affordances for model interpretability. Each semantic regex is built from primitives that span levels of abstraction, where symbols match specific characters, lexemes extend to syntactic variants, and fields encode semantic relationships. The number of components in a semantic regex is also a proxy for feature complexity, where features described using a single primitive (e.g., [:symbol left:]) are typically conceptually simpler than features that require multiple compositions of primitives and modifiers (e.g., @{:context political affiliations:}([:symbol left:]||[:symbol right:])).
We use the level of abstraction encoded in semantic regexes to measure feature complexity, finding that features become more complex deeper in the model (Figure 4). We generate semantic regexes for 1,000 features per layer in GPT-2-RES-25k, Gemma-2-2B-RES-16k, and Gemma-2-2B-RES-65k. While early-layer features are described by smaller and simpler semantic regexes, longer and more abstract semantic regexes are needed to describe later-layer features. In particular, we find that the average number of components per semantic regex (i.e., symbols, lexemes, fields, and contexts) increases across layers. This shift is mirrored in the composition of a semantic regex, where early layers have a greater proportion of single-primitive descriptions which decreases in favor of combinations of primitives and modifiers, particularly sequence and alternation compositions. Similarly, we observe that the types of primitives reflect increasing feature complexity. While all semantic regexes contain primitives, we see a decrease in low-level primitives, like lexemes, and an increase in fields (the most abstract primitive). We show an example of increasing feature complexity from early to later layers in Figure 3 bottom.
Together, these trends indicate that later layers require longer and more abstract semantic regexes to capture their complex feature behaviors. This aligns with prior research demonstrating that later layers encode increasingly complex representations (Tenney et al., 2019; Jin et al., 2025; Sun et al., 2025). However, unlike prior methods that rely on model probes or feature testing, we are able to read this complexity directly from the semantic regex feature description. As a result, while like natural language descriptions, semantic regexes allow us understand individual features, their structure also allows us to interpret model-wide attributes.
@@ 5.4 Semantic Regexes Help People Build Mental Models of LLM Features
A common role of feature descriptions is to convey the feature’s behavior to a human interpreter (Lin, 2023; Ameisen et al., 2025). Thus, we investigate how semantic regexes impact people’s mental models of LLM features (full protocol in Appendix E). We conducted a 24-person study with AI experts representative of people most likely to use feature descriptions in practice. To obtain insights across a range of features, we used 12 GPT-2-RES-25k features with diverse activation patterns and accurate natural language and semantic regex descriptions. Given a description, participants were asked to generate three activating phrases and one near-miss counterfactual.
Forming an accurate mental model of an LLM feature means being able to express the feature’s decision boundary — i.e., what does and does not activate the feature. We quantify this by measuring the difference between the feature’s maximum activation on each participant’s positive phrases and their counterfactual phrase. We choose the maximum positive phrase to account for activation artifacts that can result from tokenization. Small or negative values indicate that the participant had difficulty distinguishing activating and non-activating phrases, while large positive values reflect an understanding of the feature’s decision boundary. We compare the mean differences of max-acts and semantic-regex descriptions of the same feature, finding that participants scored as well or higher using semantic regex descriptions on 9 of 12 features (Figure 5).
@@
Although both description types were accurate, natural language often introduced extraneous details that misled participants. For instance, one feature activated on variants of the phrase “expected to”, but its natural language description included “‘expected to’ is frequently used to indicate anticipation or prediction regarding future events or outcomes”. A participant over-indexed on the additional detail, expecting the highly activating phrase “He does not know meaning of the phrase expected to” to be a counterexample because it did not indicate anticipation. In contrast, [:lexeme expect:][:symbol to:] more concisely expressed the activation pattern (see Section 5.2) and enabled participants to generate strongly activating positives and non-activating counterfactuals.
We also find that semantic regexes can reduce ambiguity by conveying activation patterns via example. For a feature that activates on times followed by “p.m.”, the semantic regex specifies this pattern directly as [:field time:][:symbol p.m.:], and as a result, every participant included “p.m.” in their positive phrases. In contrast, natural language descriptions can leave room for interpretation, where one participant misinterpreted “reference to time, specifically indicating hours in the ‘p.m.’ format” as any time in the evening and left off the critical “p.m.”. A similar pattern occurred in a feature for political conjunctions, where the semantic regex directly included [:symbol and:]|[:symbol or:], making the conjunction pattern unambiguous.
While concise semantic regexes generally benefited participants, there were cases where the verbosity of natural language was more effective. In the two features where natural language substantially outperformed semantic regexes, the semantic regexes were accurate but too minimal to elicit strongly activating phrases. One feature activated on days of the week used within full sentences. Its semantic regex [:field days of the week:] led many participants to produce single-word outputs (e.g., “Monday”) which activated slightly. Whereas, the natural language description “days of the week indicating specific events” prompted longer phrases with larger activations.
@@
6 Discussion and Limitations
Semantic regexes provide a structured syntax for describing LLM features. In doing so, they result in feature descriptions that are more consistent and concise than natural language while still matching its expressive power. However, designing this structured language introduces trade-offs that create both benefits and limitations of semantic regexes and point to several directions for future work.
Since natural language descriptions often contain irrelevant details, we designed semantic regexes to be concise. While this generally helps identify the pertinent activation pattern, it can produce overly terse descriptions. For example, [:field musician:] describes a feature that activates on famous musicians like “Taylor Swift”, incorrectly implying that “guitarist” would strongly activate the feature. Striking a balance between conciseness and expressivity across all features may involve adjusting model prompts, extending the language to encode hierarchical concepts (e.g., [:field musician.name:]), or using validation loops (Shaham et al., 2024) to mitigate ambiguity.
Additionally, although semantic regexes increase consistency, they do not enforce a unique mapping from activation pattern to description. Many valid semantic regexes can describe the same feature, even if some are less readable, like @{:context Germany:}([:symbol German:]). This non-uniqueness is not inherently problematic. In programming languages, for instance, there are many equivalent ways to implement the same function. However, languages often develop style guides that suggest the most readable syntax (van Rossum et al., 2001). Similar heuristics may benefit the semantic regex language, particularly when people (rather than models) are the intended interpreters.
To keep the semantic regex syntax minimal, we leave some components underspecified. For example, symbols match exact strings, but the semantic regex language does not specify if they also match string variations (e.g., “fruit” and “Fruit”). This simplicity avoids an overly complex vocabulary but can cause the model to make inconsistent assumptions across features (e.g., sometimes outputting [:symbol fruit:] and other times [:symbol fruit:]|[:symbol Fruit:]). Incorporating additional components, like case-insensitive flags, could reduce this ambiguity, especially as growing familiarity with semantic regexes may allow for more complex syntax.
These limitations also impact feature description evaluations. Since current metrics rely on discrete judgments of the feature description, issues like non-uniqueness and ambiguity can lead to false negatives and false positives for both semantic regexes and natural language methods. Developing more expressive metrics, such as continuous scoring schemes or readability evaluations, could provide a more comprehensive understanding of the differences between feature description methods.
Although our results indicate that semantic regexes are expressive, structured, and helpful for human interpretation, additional evaluations could deepen our understanding. For instance, large-scale repetitions of the quantitative analysis could measure pipeline stability and larger crowdsourced user studies could statistically quantify the value of semantic regexes and whether LLM-as-a-judge evaluations faithfully capture human preferences in this domain.
Semantic regexes are not a solution to polysemanticity. While composition helps capture simple polysemantic features (e.g., [:field clothing:]|[:field Ernest:]), high degrees of concept entanglement often result in incoherent descriptions, like @{:context syntax:}([:symbol by:]|[:symbol (:]|[:symbol last:]|[:symbol then:]|...|[:symbol ->:]). However, since semantic regexes are agnostic to the generation pipeline, they could easily slot into new methods for disentangling polysemantic activations (Kopf et al., 2025).
Semantic regexes require models to learn a novel language from only a brief description. As a result, we observe “grammatical” errors in generation and evaluation (e.g., misinterpreting a primitive). Although models also make mistakes on natural language descriptions, they typically misidentify the activation pattern rather than misunderstand the language. This gap raises questions for automated interpretability. People can learn structured languages and structured languages should better capture the computational roles of features, akin to how programming languages better represent algorithms. However, if models bias towards natural language, structured languages face a barrier to adoption in automated interpretability pipelines. Future work could investigate whether automated interpretability evaluations reflect the needs and abilities of human interpreters.
@@ 7 Conclusion
We introduce semantic regexes, a structured language for automatically describing LLM features. The semantic regex language is grounded in current understanding of LLM features. Each primitive and modifier is designed to reflect patterns observed in interpretability research — i.e., that features often respond to exact tokens (symbol), syntactic word forms (lexeme), and semantic categories (field), and their activations are domain-dependent (context) and co-occur with other patterns (Lin, 2023; Templeton et al., 2024). As our understanding of LLM representations grows, we expect the semantic regex language will evolve, with components added or altered to capture new feature behaviors. Just as there are many programming languages to meet different goals, future languages for interpretability may be developed with affordances suited to particular interpretability tasks, such as highlighting input vs. output features, describing particular model components (e.g., attention heads), or exposing safety-relevant features. The affordances of consistency, conciseness, and complexity that we build into semantic regexes expose a broader design space of structured languages for interpretability that can improve our collective understanding and control of LLMs.
Ethics Statement
As part of this research, we conducted a user study where we surveyed 24 participants from within our institution (Section 5.4). All participants gave informed consent and were informed they could withdraw at any time. We did not collect any identifying information, and all participant responses were anonymized. The study was approved under our institution’s internal user survey policies.
Reproducibility Statement
Implementation details for the semantic regex method, the baseline methods, and the evaluation metrics are in Appendix F, including full prompts and hyperparameters. Details to recreate our user study, including the survey instructions, are listed in Appendix E. Code is available at https://github.com/apple/ml-semantic-regex, and an interactive interface displaying our results is available at https://apple.github.io/ml-semantic-regex/.
@@ Acknowledgments
We thank our colleagues at Apple, with a special mention to Mary Beth Kery, Hadas Kotek, and Lindia Tjuatja, for giving feedback on early drafts of this work.
Appendix A LLM Usage Statement
Large language models (LLMs) are the subject of this work, and they were also used as general-purpose tools to assist with writing and coding. All research ideas, study designs, analyses, and substantive code implementations were developed by the authors.
Appendix B Additional Results
@@
Appendix C Ablation Study
Our experiments (Section 5) use GPT-4o-mini as both the explainer and evaluator model. This choice follows common practice in automated interpretability, where the same model is used to generate and evaluate feature descriptions (Bills et al., 2023; Gur-Arieh et al., 2025). Since verification is typically easier than generation, a model can reliably judge its descriptions, and the independence of model calls prevents conditioning effects.
To assess whether our findings depend on the choice of model, we conduct an ablation using a more capable model (GPT-4o) in place of GPT-4o-mini. This ablation tests whether the relative performance between semantic regexes and natural language descriptions is sensitive to model capability and evaluates the robustness of our results.
Our ablation study follows the same automated interpretability pipeline described in Section 4. All components of the pipeline remain unchanged, including feature extraction, activation formatting, and prompting strategy. The only difference is that GPT-4o is substituted for GPT-4o-mini in both the explanation and evaluation steps. We apply this setup to a randomly sampled subset of 520 features from across layers of GPT-2-RES-25k. For each feature, we generate descriptions using semantic-regex, max-acts, and token-act-pair using both GPT-4o and GPT-4o-mini, and we evaluate each description using the discrimination and generation metrics used in the main experiments. This design isolates the effect of model capability while preserving all other aspects of the experimental pipeline.
We report the results in Figures A2, A3 and A2. Across all feature description methods and evaluation metrics, the overall results remains the same regardless of whether the explainer and evaluator are GPT-4o or GPT-4o-mini. Semantic regexes continue to match the performance of natural language descriptions, and the relative differences between description types are nearly identical across the two model settings. While some metrics are slightly higher under GPT-4o (e.g., slight improvements in purity and responsiveness) these gains arise uniformly across methods and are well within the experimental noise.
Overall, the consistency between GPT-4o-mini and GPT-4o demonstrates that the comparative performance of semantic regexes is not sensitive to the choice of model. Given this robustness, along with the significantly improved efficiency and lower cost of GPT-4o-mini (Appendix D), we use GPT-4o-mini for all experiments reported in the main text.
Table A2: Across results on semantic-regex, max-acts, and token-act-pair feature descriptions, the choice of explainer and evaluator model (either GPT-4o or GPT-4o-mini) does not change the relative performance. These results validate our use of GPT-4o-mini in our main experiments. Each metric is computed on 520 randomly selected features from GPT-2-RES-25k and displayed as the mean $\pm$ the standard deviation.
Generation Discrimination
token-act-pair clarity detection fuzzing responsiveness purity
GPT-4o-mini $0.45\pm 0.37$ $0.79\pm 0.15$ $0.81\pm 0.15$ $0.83\pm 0.22$ $0.73\pm 0.26$
GPT-4o $0.59\pm 0.37$ $0.83\pm 0.13$ $0.82\pm 0.13$ $0.90\pm 0.17$ $0.84\pm 0.21$
max-acts
GPT-4o-mini $0.71\pm 0.36$ $0.86\pm 0.11$ $9\pm 0.10 0.88±0$$8\pm 0.20$ $0.81\pm 0.24$
GPT-4o $0.73\pm 0.33$ $0.86\pm 0.09$ $0.85\pm 0.09$ $0.91\pm 0.14$ $0.86\pm 0.19$
semantic-regex
GPT-4o-mini $0.69\pm 0.37$ $0.81\pm 0.13$ $0.83\pm 0.12$ $0.84\pm 0.21$ $0.68\pm 0.32$
GPT-4o $0.68\pm 0.36$ $0.82\pm 0.12$ $0.81\pm 0.12$ $0.84\pm 0.22$ $0.74\pm 0.30$
Figure A2: Metric box plots comparing different explainer and evaluator models (GPT-4o and GPT-4o-mini) across semantic-regex and natural language baseline methods (max-acts and token-act-pair). Results are shown for 520 features from GPT-2-RES-25k.
Figure A3: Metric distributions comparing different explainer and evaluator models (GPT-4o and GPT-4o-mini) across semantic-regex and natural language baseline methods. Results are shown for 520 features from GPT-2-RES-25k and visualized using kernel density estimation with bandwidth estimated via Scott’s rule.
Table A3: Cost analysis for feature description generation. Costs are computed from the number of input tokens (system prompt, few-shot examples, and feature input) and the number of generated output tokens. We report the mean $\pm$ std of input and output token counts across 130 features from GPT-2-RES-25k (10 per layer). Tokenization is performed using OpenAI’s GPT-4o model family tokenizer.
token-act-pair max-acts semantic-regex
[…]
Appendix D Cost Analysis
The cost of generating semantic-regex feature descriptions is comparable to the cost of natural language methods. Computing descriptions for all features in GPT-2-RES-25k would cost approximately \$65 using GPT-4o-mini and \$1,088 using GPT-4o. These values are on par with existing methods. Generating all token-act-pair descriptions would cost \$68, and all max-acts descriptions would cost \$54 using GPT-4o-mini.
The description cost per feature depends on the number of input tokens in the prompt, including the system prompt and few-shot examples ($T_{\text{prompt}}$) and the feature’s tokens ($T_{\text{feature}}$), and on the number of generated output tokens ($T_{\text{out}}$). It also depends on the API prices for input tokens ($P_{\text{in}}$) and output tokens ($P_{\text{out}}$).
$\text{Cost per feature}\;=\;P_{\text{in}}\left(T_{\text{prompt}}+T_{\text{feature}}\right)\;+\;P_{\text{out}}\left(T_{\text{out}}\right),$ (1)
We report the token counts and resulting costs for each feature description method in Table A3. Differences across methods are the result of formatting choices and the number of examples shown to the explainer. For instance, token-act-pair uses more verbose activation formatting than max-acts and semantic-regex, and semantic-regex uses fewer activating examples than max-acts but adds the semantic regex definition to the system prompt.
At the time of experimentation, the OpenAI API pricing [^2] https://platform.openai.com/docs/pricing is \$0.15/\$0.60 USD per million input/output tokens for GPT-4o-mini and \$2.50/\$10.00 USD per million input/output tokens for GPT-4o. These estimates assume no input caching, so the reported prices are an upper bound on the total cost of feature description generation.
Appendix E User Study Protocol
Here we provide the full protocol for our user study in Section 5.4.
Since we already evaluate semantic regex reliability across thousands of features in Section 5.1, the purpose of our user study is not to obtain another quantitative reliability estimate. Instead, our goal is to gain qualitative insight into how semantic regexes shape real-world AI experts’ interpretation of model features. To do this effectively, we recruit a representative sample of the practitioners most likely to use semantic regexes in practice and present them with a controlled but diverse set of model features. We intentionally keep the study size to 24 experts and 12 features to make the protocol feasible. Importantly, we find that these sample sizes are sufficient to draw insights about the benefits and limitations of semantic regexes.
@@ Appendix E User Study Protocol
Our target population consists of practitioners who would realistically interpret semantic regex feature descriptions in their real-world tasks, including interpretability research, AI safety, and model development. Therefore, we recruit 24 AI experts from within our organization who work with LLMs. These participants have a broad range of technical roles and backgrounds, ensuring that the study captures how semantic regexes are interpreted by the types of experts who would use them in practice.
@@
Our user study investigates how the format of a feature description influences people’s understanding. To ensure that participants interpret meaningful feature descriptions, we restricted to features from GPT-2-RES-25k with both accurate semantic regex and natural language descriptions. Because our goal is not to exhaustively evaluate all such features, but rather to study human interpretation across a representative sample, we focus on selecting a diverse subset.
We first filter to features whose semantic-regex and max-acts descriptions each achieve a detection score above 75%. From this filtered set, we manually select 12 features whose descriptions we verified to be accurate. Our selection process emphasizes feature diversity. We select features from across the model’s layers, resulting in 5 features from early layers (1-4), 4 from middle layers (5-8), and 3 from late layers (9-12). And, we chose features that represent diverse activation patterns, such that 2 activate on exact symbols, 1 on a lexeme, 3 on a field, 2 only activate in a particular context, and 4 were complex patterns that require concatenation or combination of semantic regex components.
This selection process allows us to study human interpretation across a controlled but diverse range of features without overburdening participants. We show all selected features and their descriptions in Figure A4.
@@
Participants were tasked with writing phrases that match or were a counterfactual to a given feature description. Each participant completed the task for three natural language feature descriptions (max-acts) and three semantic-regex feature descriptions. We showed each participant features of varying complexity, with half the participants seeing the feature descriptions in the left column of Figure A4 and the other half seeing feature descriptions from the right column. To avoid carryover effects, participants saw natural language and semantic regex descriptions from different features. To avoid learning effects, we randomized the order, with half the participants seeing natural language descriptions first and the other half seeing semantic regexes first. This process resulted in 425 positive and 143 counterfactual phrases (some participants did not generate all phrases).
The task was completed asynchronously and distributed via a survey link. The survey began with an introduction to the task, the semantic regex language, and an example of matching and counterfactual phrases (Figure A5). Then, participants were shown an explanation of the feature description type they would see first, followed by three feature descriptions they were asked to generate phrases for. This was repeated with the alternate feature description type and the final three explanations (Figures A6 and A7). For each feature description, they were asked to “Write 3 phrases/sentences that match the feature description” and “Write 1 phrase/sentence that is a counterfactual for the semantic regex feature description” (Figures A8 and A9).
@@
Appendix F Implementation Details
F.1 Feature Description Methods